OpenStudy (josee):

whats the derivative 2y(ln(x)) and the derivative of 3x(ln(y))

OpenStudy (amistre64):

implicits you mean?

OpenStudy (josee):

yesss

OpenStudy (amistre64):

[ 2y(ln(x))]' = 2y'(ln(x))+2y(ln(x))' = 2 y' (ln(x))+ 2y/x -2y/x = 2 y' (ln(x)) -2y/2xln(x) = y' -y/xln(x) = y'

OpenStudy (amistre64):

id say its the same wrap up for the other one as well

OpenStudy (amistre64):

but i read it wrong lol

OpenStudy (josee):

wait i got confused..

OpenStudy (amistre64):

[3x(ln(y))]' = 3x(ln(y)) + 3x(ln(y)) = 3x'(ln(y)) + 3x(ln(y))' = 3(ln(y)) + 3x/y y' -3(ln(y)) = 3x/y y' -3(ln(y))/3x = 1/y y' (ln(y))/x = 1/y y' y(ln(y))/x = y'

OpenStudy (amistre64):

which part?

OpenStudy (josee):

ok o those two sets are suppose to equal each other and im suppose to find y' but on the first one where you put +2y/x -2y/x...whered you get te -2y/x

OpenStudy (amistre64):

first set: [ 2y(ln(x))]' = 2y'(ln(x))+2y(ln(x))' ; this is just the product rule 2y derives to 2 y' ln(x) derives to 1/x x', but x'=1 here (dx/dx) = 2 y' (ln(x)) + 2y/x right?

OpenStudy (josee):

yeaa. i got 2y'lnx+2y/x

OpenStudy (amistre64):

ok, now solve for y'

OpenStudy (amistre64):

-2y/x and divide of that 2lnx

OpenStudy (josee):

wait noo. 2ylnx=3xlny lol so i look for y' out of that

OpenStudy (josee):

ok so for the right side you got what? i got 3lny+3xy'/y

OpenStudy (amistre64):

is this what your trying to derive? \[2y\ ln(x) = 3x\ ln(y)\]

OpenStudy (josee):

yessir

OpenStudy (amistre64):

lol ... youre hiding things form me :) \[\frac{d}{dx}(2y\ ln(x) = 3x\ ln(y))\] \[2y'\ ln(x)+2y\ ln(x)' = 3x'\ ln(y)+3x\ ln(y)'\] \[y'\ 2ln(x)+\frac{2y}{x} = 3ln(y)+\frac{3x}{y}\ y'\] \[y'\ 2ln(x) -\frac{3x}{y}\ y'= 3ln(y)-\frac{2y}{x}\] \[y'\ (2ln(x) -\frac{3x}{y})= 3ln(y)-\frac{2y}{x}\] \[y'\ = \frac{3ln(y)-\frac{2y}{x}}{2ln(x) -\frac{3x}{y}}\]

OpenStudy (josee):

the answers ar different.... how do you get rid of the fractions?

OpenStudy (amistre64):

algebra thru it ....

OpenStudy (josee):

cus thats what i got

OpenStudy (amistre64):

multiply top and bottom by xy

OpenStudy (amistre64):

\[y'\ = \frac{3ln(y)-\frac{2y}{x}}{2ln(x) -\frac{3x}{y}}*\frac{xy}{xy}\] \[y'\ = \frac{3xyln(y)-2yy}{2xyln(x) -3xx}\] \[y'\ = \frac{3xy\ ln(y)-2y^2}{2xy\ ln(x) -3x^2}\] maybe?

OpenStudy (amistre64):

it might help to know what the answer they say should look like

OpenStudy (josee):

cus thats what i got

OpenStudy (josee):

1) (y^2(lnx-1))/(2ylnx-3x) 2) (2y^2(lnx+1))/(2ylnx-3x) 3) (y(3xlny-2y))/(x(2ylnx-3x)) 4) (2y^2(lnx-1))/(2ylnx+3x) 5) (y(2x^2(lny)-3y))/(x(3y^2(lnx)-2x))

OpenStudy (amistre64):

\[\frac{y^2(lnx-1)}{2y\ lnx-3x}\] \[\frac{ 2y^2(lnx+1)}{(2y\ lnx-3x}\] \[ \frac{y(3xlny-2y)}{x(2ylnx-3x)}=\frac{3xy\ lny-2y^2}{2xy\ lnx-3x^2}\]

OpenStudy (amistre64):

id go 3